# JFR: A Jump Frontier Relaxation Strategy for Fast Bellman–Ford

### v1

@@ Abstract
We introduce Jump Frontier Relaxation (JFR), a correctness-preserving optimization framework that contracts active frontiers and propagates multi-hop improvements through jump-based scheduling. We provide formal proofs of convergence, monotonicity, and bounded relaxation complexity, while deliberately abstracting away implementation-specific data structures to prevent reverse engineering.
To analyze the computation–efficiency trade-off inherent to JFR, we decouple its performance into two metrics: the relaxation reduction factor $\rho_{ops}$ and the per-operation time cost factor $\rho_{TPR}$. Our analysis shows that JFR achieves net runtime gains precisely when $\rho_{ops}>\rho_{TPR}$, establishing a mathematically clear boundary for its effectiveness.
@@
We introduce JFR (Jump Frontier Relaxation with update ordering), a Bellman–Ford–based optimization framework enabling multi-hop Frontier Filtering and abstract jump propagation while strictly preserving correctness.
Unlike classical SPFA-SLF, whose performance monotonically deteriorates as graph density increases, JFR exhibits a rare non-monotonic behavior: when the number of vertices is fixed, small-scale edge increments (e.g., +5–15%) can unexpectedly accelerate the algorithm—reducing both runtime and total relaxation operations.
This edge-induced nonlinear acceleration arises because additional edges introduce shortcut structures that truncate long negative-weight propagation chains, enabling JFR’s jump mechanism to converge earlier. Such behavior has not been reported in prior BF/SPFA-family algorithms.
Across diverse graph families—including sparse, dense, and negative-weight graphs—JFR achieves substantial reductions in relaxation operations (25–99%), and remains competitive with or superior to SPFA-SLF even on ultra-large graphs (e.g., $\mathbf{N=20{,}000,E=295M}$). Occasional slowdowns relative to SPFA-SLF are bounded and structurally explainable.
The proposed Operational Efficiency (OE) metric highlights that lower relaxation counts directly translate to reduced memory traffic and computational effort, making JFR suitable for high-throughput and energy-sensitive applications.
Future work includes integrating high-performance update orderings, adaptive Frontier Filtering policies, and cache-aware layouts to further exploit JFR’s nonlinear structural sensitivity and push practical performance beyond current SPFA variants.
@@ 2.1 Frontier Sufficiency and Correctness
2.2 Amortized Analysis: Stability and Frontier Filtering
To theoretically justify the observed reduction in relaxation operations, we introduce a stability-based framework that models the effect of Frontier Filtering.
Definition 2.4 (Strict Locally Quiescent Property).
A vertex $v$ satisfies Strict $\tau$-Stability if it undergoes no strict improvement over a stability window of length $\tau$, after which it is considered inactive unless reactivated by a better update.
Assumption 2.1 (Empirical Stability).
For the graph families considered (e.g., low negative-cycle density), vertices satisfy consistency condition with a small constant $\tau$ relative to $|V|$. This is an empirically verifiable property, with statistical validation provided in Section 4.
Assumption 2.2 (Bounded Local Multi-Hop Propagation Cost).
Let $C_{\mathrm{LMH}}(t)$ denote the computational cost of the Local Multi-Hop (LMH) propagation step in iteration $t$. There exist implementation-agnostic constants $c_{1},c_{2}>0$ such that
$C_{\mathrm{LMH}}(t)\;\leq\;c_{1}\,\deg\!\big(F^{(t)}\big)\;+\;c_{2}\,N_{\mathrm{prop}}(t).$
where $\deg(F^{(t)})$ is the sum of degrees of frontier vertices, and $N_{\mathrm{prop}}(t)$ counts vertices whose distance is updated. This ensures that the overhead is bounded by the current frontier’s size and the resulting productive work.
@@ 2.2 Amortized Analysis: Stability and Frontier Filtering
Under the $\tau$-Stability property, let $s(v)$ be the number of times $v$ is active. Then $s(v)\leq 1+\lceil D_{v}/\tau\rceil$, where $D_{v}$ is the number of strict distance improvements. Consequently, the total number of edge inspections is bounded by:
$\sum_{t}|E_{F}^{(t)}|=\sum_{v\in V}s(v)\deg(v)\leq O\Big(n+m\cdot\frac{\bar{D}}{\tau}\Big),$
where $\bar{D}=\frac{1}{|V|}\sum_{v}D_{v}$ is the average updates per vertex. This bound theoretically justifies the reduction in operational complexity when $\tau$ is large and $\bar{D}$ is small.
Theorem 2.6 (Amortized Running Time).
Combining the bounded Local Multi-Hop Propagation cost (Assumption 2.2) and the amortized bound on edge inspections (Theorem 2.5), the total running time $T_{\mathrm{total}}$ satisfies:
$T_{\mathrm{total}}=\mathcal{O}\!\Big(n+\sum_{t=1}^{T}C_{\mathrm{LMH}}(t)+m\cdot\frac{\bar{D}}{\tau}\Big).$
the reduction in edge inspections by the factor $\tau$ theoretically offsets the overhead of the Local Multi-Hop Propagation step.
**The rigorous proofs for Theorems 2.5 and 2.6 are provided in Appendix A.**
2.3 Complexity Bounds (Implementation-agnostic)
@@
Under the JFR framework:
@@
    Number of successful relaxations: $O(|V|^{2})$.
@@
    Total edge inspection attempts: $O(|V||E|)$.
Hence worst-case matches Bellman–Ford, while successful relaxations admit a tighter combinatorial bound.
@@ 3.1 Algorithm Overview (Implementation-Agnostic)
    Frontier Relaxation: For every vertex $u\in F$, relax outgoing edges $(u,v)$:
@@
    Frontier Update: Construct the next frontier $F^{\prime}=\{v\mid d(v)\text{ decreased}\}$, applying $\tau$-Stability Frontier Filtering.
@@ 3.2 Relation to Recent Near-Linear Negative-Weight SSSP Results
Recent theoretical work has produced algorithms for negative-weight single-source shortest paths (SSSP) with near-linear complexity under specific assumptions (e.g., restricted graph classes, complex preprocessing steps) [9, 10, 11, 12]. JFR differs as it operates under a general directed graph with arbitrary real edge weights, focusing on improving the performance of the classic Bellman–Ford framework via practical, implementation-agnostic abstractions rather than relying on restrictive structural graph properties. The analysis provided herein connects the observable properties (Stability $\tau$ and Update Density $\bar{D}$) to the amortized complexity.
3.3 Practical Implications and Implementation Abstraction
JFR is designed to bridge the gap between classical complexity bounds and empirical optimizations. The algorithm achieves its speedup through:
@@ 3.3 Practical Implications and Implementation Abstraction
    Controlled Frontier Filtering: $\tau$-Stability provides a mechanism to prune stable, redundant nodes, as quantified by the $\frac{\bar{D}}{\tau}$ factor in the amortized bounds.
@@
    Controlled Local Multi-Hop Propagation: Updates are limited to the frontier, avoiding global scans and minimizing overhead.
@@
    Implementation Abstraction: The framework utilizes abstraction for key mechanisms (Jump Propagation, Frontier Filtering logic) to maintain theoretical generality. The specific low-level heuristics and priority structures used for optimal performance are implementation-specific and are not required for the correctness or the validity of the amortized bounds presented.
**The complete general algorithmic framework, JFR, is detailed in Appendix B.**
@@ 4.6.1 Operational Efficiency Estimation
Significantly Reduced Operational Count ($\sim 1259\times$ Order of Magnitude):
@@ 6 Conclusion and Future Work
    Normalized Work Reduction (NWR): Lower operation counts imply reduced total computational work. Prior studies show that such reductions correlate with reduced memory traffic and improved cache behavior, suggesting potential benefits in energy- or resource-constrained environments without making assumptions about physical energy models.
@@
This appendix provides the rigorous derivation of the amortized bounds stated in the main text, specifically relying on the Strict $\tau$-Stability Property (Definition 2.4).
@@ A.1 Edge-Inspection Decomposition
A.2 Proof of Theorem 2.5 (The Role of Stability)
@@
Consider a vertex $v$. Under the Strict $\tau$-Stability Property (Definition 2.4), $v$ is contractively removed from the frontier if it fails to improve strictly within a window of $\tau$ iterations. It is only re-inserted (re-activated) upon a new strict distance decrease.
This guarantees that for every block of $\tau$ iterations that $v$ is active, there must be at least one strict distance improvement $D_{v}$. Since the initial activation is separate, the total number of activation phases $s(v)$ is bounded by:
$s(v)\leq 1+\left\lceil\frac{D_{v}}{\tau}\right\rceil.$
Substituting this bound into Lemma .1:
$\sum_{t}|E_{F}^{(t)}|=\sum_{v\in V}s(v)\deg(v)\leq\sum_{v\in V}\left(1+\frac{D_{v}}{\tau}\right)\deg(v).$
@@
$\sum_{t}|E_{F}^{(t)}|\leq\sum_{v\in V}\deg(v)+\frac{1}{\tau}\sum_{v\in V}D_{v}\deg(v).$
This total inspection count is bounded by $O\left(n+m\cdot\frac{\bar{D}}{\tau}\right)$, where $\bar{D}$ incorporates the average weighted update count. ∎
@@
The total running time $T_{\text{total}}$ is the sum of the edge relaxation cost and the overhead of Jump Propagation steps across all $T$ iterations: Let $C_{\mathrm{LMH}}(t)$ denote the computational cost of the Local Multi-Hop (LMH) propagation step in iteration $t$.
@@
Substituting the amortized bound from Theorem 2.5 for the edge relaxation term:
$\sum_{t}c_{\text{relax}}\cdot|E_{F}^{(t)}|\leq O\left(n+m\cdot\frac{\bar{D}}{\tau}\right).$
The total time then satisfies: Here, $C_{\mathrm{LMH}}(t)$ denotes the computational cost of the Local Multi-Hop (LMH) propagation step in iteration $t$.
$T_{\mathrm{total}}=O\!\left(n+\sum_{t}C_{\mathrm{LMH}}(t)+m\cdot\frac{\bar{D}}{\tau}\right).$
This demonstrates that the savings factor $\frac{1}{\tau}$ directly contributes to the overall speedup. ∎
@@
1:Graph $G=(V,E,w)$, source $s$
@@
5:Initialize: Auxiliary metadata $\mathbf{aux}$ $\triangleright$ Tracks stability and update history
@@
8:  if MeetsStabilityCriterion($u$, $\mathbf{aux}$) then $\triangleright$ Checks $\tau$-stability criterion (Def. 2.4)
9:   Local Multi-Hop Propagation($u$, $\mathbf{d}$, $\pi$, $\mathcal{F}$) $\triangleright$ Performs multi-hop updates per Definition 2.1
@@
23:  if EvaluateFilteringCondition($\mathcal{F}$, $\mathbf{aux}$) then $\triangleright$ adaptive criterion based on frontier density
24:   FilterStableVertices($\mathcal{F}$, $\mathcal{Q}$) $\triangleright$ Prunes redundant/stable nodes to optimize convergence
@@
To theoretically justify Assumption 2.2 (Bounded Cost) and the Strict $\tau$-Stability Property (Definition 2.4) without loss of generality, we describe the logical control flow of the Local Multi-Hop Propagation mechanism utilized in our implementation.
@@
The Local Multi-Hop Propagation procedure is a depth-limited local relaxation process applied on the induced subgraph $G[F^{(t)}\cup N_{k}(F^{(t)})]$, where $N_{k}(F^{(t)})$ denotes the $k$-hop neighborhood of the active frontier. By limiting the propagation depth to a small constant $k$ (or a heuristic bound derived from structural indicators), the computational cost at iteration $t$ is tightly controlled by the local topology around the frontier:
$C_{\mathrm{LMH}}(t)=\mathcal{O}\Bigg(\sum_{v\in F^{(t)}}(\mathrm{deg}(v))^{k}\Bigg).$ (2)
This bounded-depth design ensures that each iteration remains efficient, regardless of the global graph size.
This structural bound ensures that the cost remains proportional to the frontier size $|F^{(t)}|$ rather than the global graph size $|E|$, satisfying Assumption 2.2.
@@
The mechanism achieves Strict $\tau$-Stability by enforcing local convergence within the depth-limited region before releasing vertices. Specifically, a vertex $v$ is only removed from the frontier (contracted) after the local Bellman-Ford process stabilizes its distance value against all paths within the $k$-hop window. Consequently, $d(v)$ cannot be improved again until a relaxation wave propagates from outside this local window, effectively guaranteeing stability for at least $\tau\approx k$ subsequent iterations.

### v2

@@ Abstract
We introduce Jump Frontier Relaxation (JFR), a correctness-preserving optimization framework that contracts active frontiers and propagates multi-hop improvements through a **$k$-bounded Local Multi-Hop ($\mathbf{k}$-LMH) strategy**. We provide formal proofs of convergence and bounded complexity, offering a constructive description of the underlying mechanisms to enable external validation.
To establish a rigorous theoretical foundation for JFR’s acceleration, we replace the empirical stability parameter with the explicit algorithm parameter $\mathbf{k}$, which bounds the depth of LMH propagation. We prove that the total number of relaxation operations is reduced to $O(n+m\cdot\frac{\overline{D}}{k})$. The overall runtime, however, is governed by a clear **cost-benefit relationship**: net acceleration is achieved only when the $\mathbf{1/k}$ reduction in relaxation operations outweighs the accumulated computational cost of the $\mathbf{k}$-LMH overhead ($\sum C_{\text{LMH}}$), establishing a mathematically sound boundary for its effectiveness.
@@
We introduce JFR, a Bellman–Ford–based optimization framework centered on $k$-bounded local multi-hop (LMH) propagation and Frontier Filtering, while strictly preserving correctness.
Unlike classical SPFA-SLF, whose performance monotonically worsens with increasing graph density, JFR exhibits a rare non-monotonic behavior: small edge increments (5–15%) can unexpectedly accelerate the algorithm by reducing runtime and relaxation operations.
This edge-induced nonlinear acceleration arises because additional edges create shortcut structures that truncate long negative-weight propagation chains, allowing LMH to converge earlier.
Across diverse graph families, JFR reduces relaxation operations by 25–99%. Its complexity is $O(n+m\,\bar{D}/k)$, where the overall speedup depends on a cost-benefit balance between the $1/k$ relaxation reduction and the accumulated $k$-LMH overhead.
The proposed Operational Efficiency (OE) metric shows that lower relaxation counts directly reduce memory traffic and computational effort, making JFR suitable for high-throughput and energy-sensitive applications.
Future work includes adaptive $k$-selection, improved update orderings, and cache-aware layouts to further exploit JFR’s nonlinear structural sensitivity and enhance performance beyond current SPFA variants.
@@ 2.1 Frontier Sufficiency and Correctness
2.2 Amortized Analysis: $k$-Bounded Propagation and Cost Tradeoff
To theoretically justify the observed reduction in relaxation operations, we introduce a constructive framework based on an explicit algorithm parameter $k$, which dictates the depth of local propagation.
Definition 2.4 ($k$-Bounded Local Multi-Hop Propagation).
The Local Multi-Hop Propagation (LMH) is defined as $k$-bounded, where $k\in\mathbb{N}$ is an explicit depth parameter. LMH ensures that local distance consistency is maintained for paths of length $\leq k$ within the active frontier’s neighborhood $N_{k}(F)$. (See Appendix Appendix A: Amortized Proof of Theorems 2.5 and 2.6 for formal details and Lemma .2 for the resulting $\tau\geq k$ stability guarantee.)
Assumption 2.1 ($k$-LMH Cost Bound).
Let $C_{\mathrm{LMH}}(t)$ denote the computational cost of the $k$-LMH propagation step in iteration $t$. We assume an upper bound proportional to $k$:
$C_{\mathrm{LMH}}(t)\;\leq\;c_{1}\cdot k\cdot\sum_{v\in N_{k}(F^{(t)})}\deg(v)\;+\;c_{2}\,N_{\mathrm{prop}}(t).$
This explicitly includes the parameter $k$ in the overhead cost, ensuring mathematical rigor.
@@ 2.2 Amortized Analysis: $k$-Bounded Propagation and Cost Tradeoff
Under the $k$-Bounded LMH property, let $s(v)$ be the number of times $v$ is active. The total number of edge inspections is bounded by:
$\sum_{t}|E_{F}^{(t)}|=\sum_{v\in V}s(v)\deg(v)\leq O\Big(n+m\cdot\frac{\bar{D}}{k}\Big),$
where $k$ is the algorithm’s depth parameter. This bound theoretically justifies the reduction in operational complexity when $k$ is large.
Theorem 2.6 (Amortized Running Time and Cost Tradeoff).
Combining the $k$-LMH cost (Assumption 2.1) and the amortized bound on edge inspections (Theorem 2.5), considering the priority queue overhead, the total running time $T_{\mathrm{total}}$ satisfies:
$T_{\mathrm{total}}=\mathcal{O}\!\Big(\left(n+m\cdot\frac{\bar{D}}{k}\right)\cdot\log|V|+\sum_{t=1}^{T}C_{\mathrm{LMH}}(t)\Big).$
Defense of the Logarithmic Factor: It is important to note that the inclusion of the logarithmic factor $\log|V|$ is a deliberate architectural choice. While this theoretically increases the per-operation cost compared to FIFO-based approaches ($O(1)$), the strict ordering enforced by the priority queue drastically suppresses the relaxation count term ($m\cdot\frac{\bar{D}}{k}$). The overall speedup is achieved when the reduction in total relaxations outweighs the accumulated overhead.
The rigorous proofs for Theorems 2.5 and 2.6, including the detailed derivation of the $k$-dependent bounds, are provided in Appendix A.
2.3 Complexity Bounds and Robustness
@@
Under the JFR framework using a priority-based implementation:
@@
    Worst-case time complexity: $O(|V||E|\log|V|)$.
@@
    Robustness Advantage: Unlike SPFA-SLF, whose queue-based dynamics are known to exhibit extremely poor worst-case behavior—with repeated oscillations causing up to $\Theta(|V||E|)$ relaxations—the JFR framework maintains a strictly polynomial upper bound. While certain adversarial inputs make SPFA-SLF appear to grow “faster than polynomial” in practice, its formal worst-case time complexity remains bounded by $O(|V||E|)$.
Hence, JFR sacrifices a logarithmic factor to ensure robustness against the exponential degradation observed in heuristic variants.
Note: While the worst-case bound includes $\log|V|$, the practical efficiency is captured by the much tighter amortized edge inspection bound of $O(n+m\cdot\frac{\bar{D}}{k})$, as rigorously shown in Theorem 2.5.
@@ 3.1 Algorithm Overview (Implementation-Agnostic)
    Frontier Relaxation. For every vertex $u\in F$, relax all outgoing edges $(u,v)$.
@@
    Frontier Update: Construct the next frontier $F^{\prime}=\{v\mid d(v)\text{ decreased}\}$, leveraging the stability provided by the $\mathbf{k}$-bounded LMH to prune stable nodes.
@@ 3.2 Relation to Recent Near-Linear Negative-Weight SSSP Results
Recent theoretical work has produced algorithms for negative-weight single-source shortest paths (SSSP) with near-linear complexity under specific assumptions (e.g., restricted graph classes, complex preprocessing steps) [9, 10, 11, 12]. JFR differs as it operates under a general directed graph with arbitrary real edge weights, focusing on improving the performance of the classic Bellman–Ford framework via practical, implementation-agnostic abstractions rather than relying on restrictive structural graph properties. The analysis provided herein connects the observable properties (Stability and Update Density $\bar{D}$) to the amortized complexity.
3.3 Practical Implications and Implementation Generality
JFR is designed to bridge the gap between classical complexity bounds and empirical optimizations. The algorithm achieves its speedup through mechanisms now formally linked to the parameter $k$:
@@ 3.3 Practical Implications and Implementation Generality
    Controlled Frontier Filtering. The stability property (i.e., $\tau\geq k$) provides a mechanism to prune stable or redundant nodes. The resulting speedup is reflected in the amortized factor $\bar{D}/k$, provided that the associated overhead remains low.
@@
    Controlled Local Multi-Hop Propagation. Updates are limited to the $k$-hop neighborhood of the frontier, avoiding global scans and minimizing overhead.
@@
    Implicit $k$-Boundedness via Priority Queue. In our high-performance implementation, the theoretical concept of $k$-bounded LMH is realized implicitly via a Priority Queue (PQ). The PQ dynamically determines the effective propagation depth by always prioritizing the most promising nodes (the ’Jump’). This effectively filters out suboptimal, shallow updates that would otherwise clutter the frontier, creating a dynamic, self-adjusting $k$ that adapts to the local graph topology without manual tuning.
  • 4.
    Implementation Generality. The framework utilizes general abstractions for key mechanisms (Jump Propagation, Frontier Filtering logic) to maintain theoretical generality. While our reference implementation relies on a priority structure to achieve the dynamic $k$-bound, other low-level heuristics (such as multi-level buckets) could also satisfy the amortized bounds presented.
The complete general algorithmic framework, JFR, is detailed in Appendix B.
@@ 4.6.1 Operational Efficiency Estimation
Significantly Reduced Operational Count ($\sim 3\times$ Order of Magnitude):
@@ 6 Conclusion and Future Work
    Normalized Work Reduction (NWR). Lower operation counts imply a reduction in total computational work. Prior studies show that such reductions correlate with lower memory traffic and improved cache behavior. This connection has also been observed in empirical energy studies of shortest-path algorithms; for example, Alamoudi and Al-Hashimi [16] report that reduced operation counts in Bellman–Ford variants directly translate to smoother memory-access patterns and more energy-efficient execution. This suggests that JFR’s work reductions may offer benefits in energy- or resource-constrained environments, even without assuming a specific physical energy model.
@@
This appendix provides the rigorous derivation of the amortized bounds stated in the main text, specifically relying on the $k$-Bounded Local Multi-Hop Property (Definition 2.4).
@@ A.1 Edge-Inspection Decomposition
A.2 Proof of Theorem 2.5 (The Role of $k$-Bounded Propagation)
Lemma .2 ($k\Rightarrow\tau\geq k$).
If LMH is $k$-bounded and a vertex $v$ is stabilized by LMH in iteration $t$ (i.e., no strict improvement is available via $\leq k$-length paths inside $N_{k}(F(t))$), then any strict distance improvement to $v$ originating from outside the $k$-hop neighborhood $N_{k}(F(t))$ requires at least $k$ outer iterations to propagate to $v$. Consequently, the observed stability window $\tau$ satisfies $\tau\geq k$.
@@
Any external strict improvement must traverse at least one edge to enter the $k$-hop neighborhood $N_{k}(F(t))$. Since $k$-bounded LMH exhaustively stabilizes all $\leq k$-length paths within this region, the propagation of the external effect must proceed via outer-iteration frontier expansions, each of which advances the affected region by at most one hop. Thus, at least $k$ outer iterations are required before the external update can reach $v$. ∎
Proof of Theorem 2.5.
Consider a vertex $v$. Under the $k$-Bounded Local Multi-Hop Property (Definition 2.4) and the resulting stability lower bound $\tau\geq k$ (Lemma .2), a vertex $v$ can be strictly improved at most once every $k$ outer iterations. It is only re-inserted (re-activated) upon a new strict distance decrease.
This guarantees that for every block of $k$ outer iterations, there must be at most one strict distance improvement $D_{v}$ that activates $v$. Since the initial activation is separate, the total number of activation phases $s(v)$ is bounded by:
$s(v)\leq 1+\left\lceil\frac{D_{v}}{k}\right\rceil.$
Substituting this new bound (using the algorithm parameter $k$ instead of the observable $\tau$) into Lemma .1:
$\sum_{t}|E_{F}^{(t)}|=\sum_{v\in V}s(v)\deg(v)\leq\sum_{v\in V}\left(1+\frac{D_{v}}{k}\right)\deg(v).$
@@
$\sum_{t}|E_{F}^{(t)}|\leq\sum_{v\in V}\deg(v)+\frac{1}{k}\sum_{v\in V}D_{v}\deg(v).$
This total inspection count is bounded by $O\left(n+m\cdot\frac{\bar{D}}{k}\right)$. ∎
@@
The total running time $T_{\text{total}}$ is the sum of the edge relaxation cost and the total overhead of $k$-Bounded Local Multi-Hop (LMH) Propagation steps across all outer iterations $T$. Let $C_{\mathrm{LMH}}(t)$ denote the computational cost of the LMH propagation step in iteration $t$, which is bounded by Assumption 2.1.
@@
Substituting the edge inspection bound from Theorem 2.5 for the edge relaxation term:
$\sum_{t}c_{\text{relax}}\cdot|E_{F}^{(t)}|\leq O\left(n+m\cdot\frac{\bar{D}}{k}\right).$
The total time then satisfies:
$T_{\mathrm{total}}=O\!\left(  n  +m\cdot\frac{\bar{D}}{k}  +\sum_{t}C_{\mathrm{LMH}}(t)\right).$
This revised bound explicitly shows that the overall speedup depends on two competing factors: the savings factor $1/k$ applied to the classical relaxation term, and the accumulated overhead cost $\sum_{t}C_{\mathrm{LMH}}(t)$, which is proportional to $k$ (Assumption 2.1). For a net speedup, the benefit from the $1/k$ reduction must outweigh the cost term. ∎
@@
1:Graph $G=(V,E,w)$, source $s$, depth parameter $k$
@@
5:Initialize: Auxiliary metadata $\mathbf{aux}$ $\triangleright$ Tracks update history and local stability
@@
8:  if MeetsStabilityCriterion($u$, $\mathbf{aux}$) then $\triangleright$ Checks if node is locally stable (implies $\tau\geq k$, see Lemma .2)
9:   Local Multi-Hop Propagation($u$, $\mathbf{d}$, $\pi$, $\mathcal{F},k$) $\triangleright$ Performs $k$-bounded updates per Definition 2.4
@@
23:  if EvaluateFilteringCondition($\mathcal{F}$, $\mathbf{aux}$) then $\triangleright$ Adaptive criterion based on frontier density and $k$
24:   FilterStableVertices($\mathcal{F}$, $\mathcal{Q}$) $\triangleright$ Prunes redundant nodes leveraging $k$-stability
@@
To theoretically justify Assumption 2.1 ($k$-LMH Cost Bound) and the mechanisms described in Definition 2.4 ($k$-Bounded Local Multi-Hop Propagation) without loss of generality, we describe the logical control flow of the Local Multi-Hop Propagation mechanism utilized in our implementation.
@@
The Local Multi-Hop Propagation procedure is a depth-limited local relaxation process applied on the induced subgraph $G[N_{k}(F^{(t)})]$, where $N_{k}(F^{(t)})$ denotes the $k$-hop neighborhood of the active frontier. By limiting the propagation depth to a small constant $k$ (or a heuristic bound derived from structural indicators), the computational cost at iteration $t$ is tightly controlled by the local topology around the frontier. Consistent with Assumption 2.1, the cost is bounded by the size of the local neighborhood and the depth parameter $k$:
$C_{\mathrm{LMH}}(t)=\mathcal{O}\Bigg(k\cdot\sum_{v\in N_{k}(F^{(t)})}\mathrm{deg}(v)\Bigg).$ (2)
This bounded-depth design ensures that each iteration remains efficient. The cost remains proportional to the local frontier neighborhood size rather than the global graph size $|E|$, satisfying the conditions for the cost-benefit tradeoff derived in Theorem 2.6.
@@
The mechanism achieves the stability guarantees proved in Lemma .2 ($\tau\geq k$) by enforcing local convergence within the depth-limited region before releasing vertices. Specifically, a vertex $v$ is only removed from the frontier (contracted) after the local relaxation process stabilizes its distance value against all paths of length $\leq k$ within the window. Consequently, $d(v)$ cannot be improved again until a relaxation wave propagates from outside this local window, effectively guaranteeing stability for at least $\tau\geq k$ subsequent outer iterations.